OpenStudy (anonymous):

What is the equation of a line that passes through the point (-2, 5) and is parallel to a line with a slope of 3 so it is like y=3x+1

OpenStudy (anonymous):

slope is always x

OpenStudy (aum):

slope is m = 3 ( not b = 3)

OpenStudy (aum):

y = mx + b slope m = 3 y = 3x + b passes through (-2,5). Put x = -2 and y = 5 5 = 3(-2) + b 5 = -6 + b b = 11 y = 3x + 11

OpenStudy (anonymous):

y = 3x + 1 y = 3x + 5 y = 3x + 7 y = 3x + 11 these are the answers the letter before x is always the slope and i plugged in the points into the equation and got A is that right

OpenStudy (anonymous):

got it i subtracted instead of added your a lifesaver

OpenStudy (anonymous):

Wow at first glance i thought your name was destiny i shot

OpenStudy (aum):

You are welcome.

OpenStudy (anonymous):

haahahahha they had taken all of the names so im not self absorbed and yes i saw that afterward and laughed to......

OpenStudy (anonymous):

What is the equation of a line that passes through the point (1, 3) and is parallel to a line with a slope of -2

OpenStudy (anonymous):

so y=-2x=5

OpenStudy (aum):

y = -2x + 5